OpenStudy (anonymous):

Find the sum of the series 1+ln(2)+(ln 2)^2/2!+...+(ln 2)^n/n!+...?

OpenStudy (kainui):

Well since e^x=1+x+x^2/2!+...+x^n/n! then you're really doing e^ln2 which just equals 2.

OpenStudy (anonymous):

So for problems like these, you must first memorize the Maclaurin series, right?

OpenStudy (kainui):

I suppose so. Often times you can recognize a Maclaurin series based on how its derivatives behave. For instance, e^x is it's own derivative right? Similarly, that polynomial 1+x+x^2/2!+...+x^n/n!+... is also its own derivative.

OpenStudy (kainui):

Which makes perfect sense, since they're exactly equal!

OpenStudy (anonymous):

Thank you!

OpenStudy (kainui):

One way to remember sine and cosine is by remembering that: e^(ix)=cosx+i*sinx So by having e^x memorized, simply plugging in (ix) for (x) will give you the series for cosine (which is all even polynomial terms, 1, x^2, x^4, etc...) and the i will make it alternating.